OpenStudy (anonymous):

Factor the expression: x3 - 2x2 - 9x + 18

OpenStudy (anonymous):

You have to be really careful when there's a negative sign between the second and third term. x^3+2x^2-9x-18 = x^2(x+2) - 9(x+2) Notice that when you distribute the -9, the + will change to a -. You can simplify further: (x^2-9)(x+2) = (x+3)(x-3)(x+2)

OpenStudy (anonymous):

Do u understand now?

OpenStudy (anonymous):

on last part is it x+2 or x-2?

OpenStudy (anonymous):

+

OpenStudy (anonymous):

it doesnt have that as a option

OpenStudy (anonymous):

what are your options?

OpenStudy (anonymous):

(x - 3)(x - 3)(x - 2) (x + 3)(x + 3)(x - 2) (x + 3)(x - 3)(x - 2) (x + 3)(x + 3)(x +2)

OpenStudy (anonymous):

its the last one like i said

OpenStudy (anonymous):

on urs middle is x-3

OpenStudy (anonymous):

well thats what i got but give the last one a try

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

your welcome :)